OpenStudy (thomasfire55):

Divide the following polynomial using synthetic division, then place the answer in the proper location on the grid. Write answer in descending powers of x. (x^4 -81)/(x-3)

jhonyy9 (jhonyy9):

for x^4 -81 you can using formula of difference of two perfect squares do you know it ?

jhonyy9 (jhonyy9):

or you need rewriting the x^4 -81 = x^4 +0x^3 +0x^2 +0x -81 / x - 3 = yes ?

jhonyy9 (jhonyy9):

@TheSmartOne what you think please about this ?

jhonyy9 (jhonyy9):

ok. ty

TheSmartOne (thesmartone):

x^4 - 81 Use a^2 - b^2 = (a+b)(a-b) x^4 - 81 = (x^2)^2 - (9)^2

jhonyy9 (jhonyy9):

using synthetic division - sorry but i not heard about nothing

TheSmartOne (thesmartone):

http://www.purplemath.com/modules/synthdiv.htm

jhonyy9 (jhonyy9):

thank you very much - this i know but not named synth divid.